# Investigating the role of mutual information in the Page curve for a functional renormalization group improved Schwarzschild black hole

### v1

@@ II After Page time scenario and the role of mutual information of subsystems in the Page curve
As we have already discussed, the paradoxical situation arises just after the Page time when fine-grained entropy of Hawking radiation exceeds the thermal (coarse-grained) entropy of the black hole and the inclusion of the island contributions helps us to overcome this unease situation. In short, we need to make use of the formula given in eq.(1) in order to compute the correct form of fine-granied entropy of black hole radiation. The second term in eq.(1), can be computed by considering the fact that, the quantum state on the full Cauchy slice is a pure state. This implies one can write down the following equality regarding the relevant von-Neumann entropies
@@
The regions $B_{\pm}$ are are extended from the points $b_{\pm}=(\pm t_{b},b)$ to the boundaries of the island region (location of the quantum extremal surfaces), that is, $a_{\pm}=(\pm t_{a},a)$. This can be graphically represented with the help of a Penrose diagram, as given in Figure (2). In order to compute the von-Neumann entropy of the matter fields, we have considered that, the entire spacetime is filled with $2d$ free conformal matter (with large central charge $c$).
@@
In the present scenario, our interest lies in the computation of the von-Neumann entropy of matter fields on two disjoint subsystem, $S_{vN}(B_{+}\cup B_{-})$. This we can do by using the following formula Calabrese et al. (2009)
@@
One then needs to substitute the above in eq.(1) and perform the extremization of the island parameters $t_{a}$ and $a$ which results in $t_{a}\approx t_{b}$ and $a\approx r_{+}$. Now, substition of these extremized values leads one to the following form of fine-grained entropy of radiation after the Page time
@@
$a$ $=$ $r_{+}+\left(\frac{cG}{24\pi}\right)\frac{1}{r_{+}}+...$ (37)
$=$ $a^{(S)}+\frac{\tilde{\omega}}{M}\left(\frac{Gc}{48\pi}\frac{1}{(r^{(S)})^{2}}-\frac{1}{2}\right)$
where $a^{(S)}$ is the extremized value of ``$a"$ which one can obtain for a usual classical Schwarzschild Schwarzshild black hole, given in Saha et al. (2022) 444 The expression of $a_{(S)}$ given by, $a^{(S)}=r^{(S)}+\left(\frac{cG}{24\pi}\right)\frac{1}{r^{(S)}}$.. It is to be mentioned that $r^{(S)}$ is the horizon radius of Schwarzschild black hole. The second term in the above result arises due to the inclusion of the quantum corrections to the classical geometry. The expression for $r_{+}$ has been provided in eq.(14). In the limit $\tilde{\omega}\rightarrow 0$, one can recover the result of island parameter for Schwarzschild black hole.
@@
$S(R)$$\approx$ $2S_{BH}^{(q)}-\frac{c}{4}\log\left(S_{BH}^{(q)}\right)+\frac{\left(\frac{c}{2}\right)^{2}}{2S_{BH}^{(q)}}+...\leavevmode\nobreak$ (38)
@@
$S(R)$ $=$ $S^{(S)}(R)+\Bigg{[}\frac{\frac{c}{3}}{2S_{BH}^{(S)}}\left(\frac{\tilde{\omega}}{2M^{2}G}\right)-2S_{BH}^{(S)}\left(\frac{\tilde{\omega}}{4M^{2}G}\right)$ (39)
$+$ $\frac{c}{3}\left(\frac{\tilde{\omega}}{4M^{2}G}\right)\left(1+\frac{\tilde{\omega}}{8M^{2}G}\right)\Bigg{]}$
where $S^{(S)}(R)$ denotes the fine-grained entropy of Hawking radiation for the standard, classical Schwarzschild black hole solution Saha et al. (2022) [^5] The expression of $S^{(S)}(R)$ reads, $S^{(S)}(R)=2S^{(S)}_{BH}+\frac{c}{4}\log S^{(S)}_{BH}+\frac{(c/3)^{2}}{2S^{(S)}_{BH}}$..
@@
$t_{Scr}^{(q)}$ $=$ $t_{Scr}^{(S)}+\frac{\beta^{(S)}}{4\pi}\Bigg{[}\frac{\tilde{\omega}^{2}}{4G^{2}M^{4}}\log\left(S_{BH}^{(S)}\left(1-\frac{\tilde{\omega}}{2r^{(S)}M}\right)^{2}\right)$ (41)
$+$ $2\log\left(1-\frac{\tilde{\omega}}{2r^{(S)}M}\right)\Bigg{]}$
where $t_{Scr}^{(S)}$ is the scrambling time for the classical Schwarzschild black hole  [^6] The scrambling time for Schwarzschild black hole reads $t_{Scr}^{(S)}=\frac{\beta^{(S)}}{4\pi}\log S_{BH}^{(S)}$. To get the the above result we have used the relations given in eq.(14) and eq.(18). It is interesting to observe that the above result suggests $t_{Scr}^{(q)}-t_{Scr}^{(S)}>0$, this in turn implies that the information recovery time is larger for the QCSBH in comparison to the standard Schwarzschild black hole.
Now we will make few comments regarding the obtained result. Keeping in mind the relation between entanglement wedge and mutual information Takayanagi and Umemoto (2018), one can argue that as long as $t_{a}-t_{b}<t_{Scr}^{(q)}$, the entanglement wedge associated with $B_{+}\cup B_{-}$ is in the connected phase as the mutual information between $B_{+}$ and $B_{-}$ is nonzero. However, as soon On the other hand, as soos as the difference between $t_{a}-t_{b}$ equals $t_{Scr}^{(q)}$, the entanglement wedge associated with $B_{+}\cup B_{-}$ is in the disconnected phase as mutual information between $B_{+}$ and $B_{-}$ vanishes.
Finally, we will compute the Page time $(t_{P}^{(q)})$ for the QCSBH. This can be obtained by equating the leading order result of eq.(51) and eq.(38). The expression of $(t_{P}^{(q)})$ reads
@@
$t_{P}^{(q)}\approx t_{P}^{(S)}\left(1-\frac{\tilde{\omega}}{r^{(S)}M}\right)$ (43)
where $t_{P}^{(S)}$ is the Page time for the standard Schwarzschild black hole, the expression of $t_{P}^{(S)}$ reads $t_{P}^{(S)}=\frac{3\beta^{S}}{\pi c}S_{BH}^{S}$. As we have aleady mention that the island starts contribute to the von-Neumann entropy of Hawking radiation just after the Page time. Therefore, our result suggests that for a quantum-corrected black hole, island formation begins earlier compared to the Schwarzschild black hole.
@@
$=$ $\textrm{min}\leavevmode\nobreak\ \mathop{\textrm{ext}}_{\mathcal{\mathrm{a}}}\bigg{\{}\frac{\textrm{Area}(\partial I)}{4G_{N}}+S_{vN}(B_{+})+S_{vN}(B_{-})\bigg{\}}\leavevmode\nobreak\ .$
In the first line we make use of the fact that the state on the full Cauchy slice is a pure state and based upon this observation one can write $S_{vN}(B_{+}\cup B_{-})=S_{vN}(I\cup R)$. Furthermore, it is subject to the condition of vanishing of mutual information between $B_{+}$ and $B_{-}$. In the final form of the formula we have incorporated the the fact $t_{a}-t_{b}=|r^{*}(a)*r^{*}(b)|$ which in turn fixes one of the island parameters, namely, $t_{a}$ and only ``$a"$ is left to be extremized. We propose the above formula for evaluating the fine-grained entropy of Hawking radiation for an eternal black hole.
@@ III Before Page time scenario and the emergence of Hartman-Maldacena time
$S(R)$$\approx$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{\pi}\right)\sqrt{f(b)}\right]+\left(\frac{c}{6}\right)\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)^{2}\leavevmode\nobreak\ .$ (50)
@@
$S(R)$$\approx$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{\pi}\right)\sqrt{f(b)}\right]+\left(\frac{c}{3}\right)\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right).\leavevmode\nobreak\ \leavevmode\nobreak$ (51)
From the above given expressions one can note that in the early time domain $S(R)$ has quadratic time dependence, that is $S(R)\sim t_{b}^{2}$. On the other hand in the late time domain, $S(R)$ grows linearly over observer’s time, $S(R)\sim t_{b}$ which matches with Hawking’s suggestion. This observation is agrees with Hartman and Maldacena (2013).
In this section out aim is to compute the mutual information between the matter fields localized on region $R_{+}$ and $R_{-}$. The reason behind this interest is the following. In the last section, we have already seen the importance of mutual information in obtaining the Page curve. In the before Page time scenario, although there is no island contribution but it will be very interesting to see how does the mutual information evolves over time. To compute the von Neumann entropy of matter fields on $R_{\pm}$, we will use the formula given below Calabrese and Cardy (2009)
@@
$d(b_{\pm},e_{\pm})$
$=$ $\sqrt{2F(b)F(e)e^{\kappa^{(q)}r^{*}(b)}[\cosh(\kappa^{(q)}r^{*}(b))-\cosh(\kappa^{(q)}t_{b})]}\leavevmode\nobreak\ .$
@@
$S_{vN}(R_{+})=S_{vN}(R_{-})=\left(\frac{c}{6}\right)\log\left[2\left(\frac{\beta^{(q)}}{\pi}\right)^{2}\sqrt{f(b)f(e)}\left\{|\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)-\cosh\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)|\right\}\right]\leavevmode\nobreak\ .$ (54)
@@
$=$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(b)}\left\{\frac{|\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)-\cosh\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)|}{\cosh\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)}\right\}\right]\leavevmode\nobreak\ .$
For the sake of better understanding, we look at the behaviour of $I(R_{+}\cup R_{-})$ in two different time domains, namely in the early time and in the late time domain. In the early time domain, that is when $t_{b}\ll\beta^{(q)}$ the expression of the mutual information between $R_{+}$ and $R_{-}$ reduces to the following form
$I(R_{+}:R_{-})$$\approx$ $\left(\frac{c}{3}\right)\left[\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(\xi)}\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right]-\mathrm{sech}\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)-\left(\frac{2\pi^{2}}{\beta^{(q)^{2}}}\right)\left\{1+\mathrm{sech}\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right\}t_{b}^{2}\right]\leavevmode\nobreak\ .$
The above expression suggests that at the very begining, that is, at $t_{b}=0$, $I(R_{+}:R_{-})$ is non-zero and finite. Further, it then decreases with respect to the observer’s time $t_{b}$. On the other hand at late time, that is for $t_{b}\gg\beta^{(q)}$, $I(R_{+}:R_{-})$ boils down to the following form
$I(R_{+}:R_{-})$$\approx$ $\left(\frac{c}{3}\right)\left[\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(\xi)}\right]-2\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)e^{-\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)}\right]\leavevmode\nobreak\ .$ (57)
The above result of $I(R_{+}:R_{-})$ implies that, at late time $I(R_{+}:R_{-})$ increases with respect to $t_{b}$. This implies the behaviour of $I(R_{+}:R_{-})$ in two different time domains are different, in particular at the early time $I(R_{+}:R_{-})$ is a decreasing function with respect to time and in the late time domain it is increasing. This observation motivates us to look for a time scale $t_{R}$ where the mutual information between $R_{+}$ and $R_{-}$ vanishes as zero is the lowest possible value for non-negative quantity like $I(R_{+}:R_{-})$. This in turns means that when $t_{b}=t_{R}$ the entanglement wedge associated with $R_{+}\cup R_{-}$ become disconnected. This was first addressed in Roy Chowdhury et al. (2022, 2023) with the help of the following proposal
@@
$t_{R}=$ $\left(\frac{\beta^{(q)}}{2\pi}\right)\times$
$\cosh^{-1}\left\{\left(\frac{\frac{\beta^{(q)}}{2\pi}\sqrt{f(e)}}{1+\frac{\beta^{(q)}}{2\pi}\sqrt{f(e)}}\right)\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right\}.$
@@
$S_{vN}^{t_{b}=t_{R}}(R_{+}\cup R_{-})$ $=$ $\frac{c}{3}\log\left[\frac{\left(\frac{\beta^{(q)}\sqrt{f(e)}}{2\pi}\right)^{2}}{1+\frac{\beta^{(q)}\sqrt{f(\xi)}}{2\pi}}\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right]$
@@
So, our proposal suggests that in the time domain $0\leq t_{b}<t_{R}$, the mutual information between the matter fields on $R_{+}$ and $R_{-}$ is non zero and the entanglement wedge associated with $R_{+}\cup R_{-}$ is in its connected pahse. However, at $t_{b}=t_{R}$ the mutual correlation vanishes ($I(R_{+}\cup R_{-})=0$), this means that at $t_{b}=t_{R}$ the entanglement wedge corresponding to $R_{+}\cup R_{-}$ is in the disconnected pahse. These observations strongly indicate that this time-scale $t_{R}$ is nothing but the Hartman-Maldacena time $t_{HM}$. This observation agrees with the one given in Grimaldi et al. (2022). Further, the expression of $I(R_{+}:R_{-})$ at $t_{b}=\beta^{(q)}$ (by substituting $t_{b}=\beta^{(q)}$ in eq.(III)), is given by
$I(R_{+}:R_{-})$ $=$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(\xi)}\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right]\leavevmode\nobreak\ .$

### v2

@@ II After Page time scenario and the role of mutual information of subsystems in the Page curve
As we have already discussed, the paradoxical situation arises just after the Page time when fine-grained entropy of Hawking radiation exceeds the thermal (coarse-grained) entropy of the black hole and the inclusion of the island contributions helps us to overcome this unease situation. In short, we need to make use of the formula given in eq.(1) in order to compute the correct form of fine-granied entropy of black hole radiation. The second term in eq.(1), can be computed by considering the fact that, the quantum state on the full Cauchy slice is a pure state. This implies one can write down the following equality regarding the relevant von Neumann entropies
@@
The regions $B_{\pm}$ are are extended from the points $b_{\pm}=(\pm t_{b},b)$ to the boundaries of the island region (location of the quantum extremal surfaces), that is, $a_{\pm}=(\pm t_{a},a)$. This can be graphically represented with the help of a Penrose diagram, as given in Figure (2). In order to compute the von Neumann entropy of the matter fields, we have considered that, the entire spacetime is filled with $2d$ free conformal matter (with large central charge $c$).
@@
In the present scenario, our interest lies in the computation of the von Neumann entropy of matter fields on two disjoint subsystem, $S_{vN}(B_{+}\cup B_{-})$. This we can do by using the following formula Calabrese et al. (2009)
@@
One then needs to substitute the above in eq.(1) and perform the extremization with respect to the island parameters $t_{a}$ and $a$, which results in $t_{a}\approx t_{b}$ and $a\approx r_{+}$. Now, substitution of these extremized values of island parameters, leads to the following form of fine-grained entropy of radiation after the Page time
@@
This proposal departs from the standard island prescription where we find that the mutual information between the regions $B_{+}$ and $B_{-}$ vanishes approximately (eq.(II)). Hence, in the standard island prescription, one requires extremization of the fine grained entropy with respect to the island parameters. As we shall find in the subsequent discussion, the vanishing of the mutual information between the regions $B_{+}$ and $B_{-}$ fixes the island parameter $t_{a}$.
@@
$a$ $=$ $r_{+}-\left(\frac{cG_{0}}{24\pi}\right)\frac{1}{r_{+}}+...$ (38)
$\approx$ $a^{(S)}-\frac{\tilde{\omega}}{M}\left(\frac{cG_{0}}{48\pi}\frac{1}{r^{(S)}}+\frac{1}{2}\right)$
where $a^{(S)}$ is the extremized value of ``$a"$ which one can obtain for a usual classical Schwarzschild Schwarzshild black hole, given in Saha et al. (2022) 444 The expression of $a_{(S)}$ given by, $a^{(S)}=r^{(S)}-\left(\frac{cG}{24\pi}\right)\frac{1}{r^{(S)}}$.. It is to be mentioned that $r^{(S)}$ is the horizon radius of Schwarzschild black hole. The second term in the above result arises due to the inclusion of the quantum corrections to the classical geometry. The expression for $r_{+}$ has been provided in eq.(14). In the limit $\tilde{\omega}\rightarrow 0$, one can recover the result of island parameter for standard Schwarzschild black hole.
@@
$S(R)$$\approx$ $2S_{BH}^{(q)}-\frac{c}{6}\log\left(S_{BH}^{(q)}\right)-\frac{\left(\frac{c}{12}\right)^{2}}{2S_{BH}^{(q)}}+...\leavevmode\nobreak$ (39)
@@
$S(R)$$\approx$ $S^{(S)}(R)+\tilde{\omega}\Bigg{[}\frac{\frac{c}{3}}{4G_{0}M^{2}}-\frac{S_{BH}^{(S)}}{M^{2}G_{0}}-\frac{\left(\frac{c}{12}\right)^{2}}{S_{BH}^{(S)}}\frac{1}{4M^{2}G_{0}}\Bigg{]}...$ (40)
$=$ $S^{(S)}(R)+\tilde{\omega}\pi\Bigg{[}\frac{\frac{c}{3}}{S_{BH}^{(S)}}-\frac{\left(\frac{c}{12}\right)^{2}}{(S_{BH}^{(S)})^{2}}-4\Bigg{]}$
where $S^{(S)}(R)$ denotes the fine-grained entropy of Hawking radiation for the standard, classical Schwarzschild black hole solution Saha et al. (2022) [^5] The expression of $S^{(S)}(R)$ reads, $S^{(S)}(R)=2S^{(S)}_{BH}-\frac{c}{6}\log S^{(S)}_{BH}-\frac{(c/12)^{2}}{2S^{(S)}_{BH}}$..
@@
$t_{Scr}^{(q)}$$\approx$ $t_{Scr}^{(S)}+\frac{\beta^{(S)}}{4\pi}\Bigg{[}\frac{\tilde{\omega}G_{0}}{(r^{(S)})^{2}}\log S_{BH}^{(S)}$ (42)
$+$ $\left(1+\frac{\tilde{\omega}G_{0}}{(r^{(S)})^{2}}\right)\log\left(1-\frac{\tilde{\omega}}{Mr^{(S)}}\right)\Bigg{]}+...$
$=$ $t_{Scr}^{(S)}+\frac{\beta^{(S)}}{4\pi}\Bigg{[}\frac{\pi\tilde{\omega}}{S_{BH}^{(S)}}\log S_{BH}^{(S)}$
$+$ $\left(1+\frac{\pi\tilde{\omega}}{S_{BH}^{(S)}}\right)\log\left(1-\frac{\tilde{2\pi\omega}}{S_{BH}^{(S)}}\right)\Bigg{]}$
where $t_{Scr}^{(S)}$ is the scrambling time for the classical Schwarzschild black hole  [^6] The scrambling time for Schwarzschild black hole reads $t_{Scr}^{(S)}=\frac{\beta^{(S)}}{4\pi}\log S_{BH}^{(S)}$. To get the the above result we have used the relations given in eq.(14) and eq.(18). It is interesting to observe that the above result suggests $t_{Scr}^{(q)}-t_{Scr}^{(S)}>0$, this in turn implies that the information recovery time is larger for the QCSBH in comparison to the standard Schwarzschild black hole. The fact that the information recovery time is longer in case of the QCSBH refers to the fact that it takes more time to scramble the information in case of the QCSBH which arises from the asymptotic safe theory of quantum gravity. This is in line with the fact that in a quantum system, the scrambling time is longer. Hence, the entire process of recovering the information after it has been scrambled in the QCSBH is expected to be longer compared to the standard Schwarzschild black hole.
Now we will make few comments regarding the obtained result. Keeping in mind the relation between entanglement wedge and mutual information Takayanagi and Umemoto (2018), one can argue that as long as $t_{a}-t_{b}<t_{Scr}^{(q)}$, the entanglement wedge associated with $B_{+}\cup B_{-}$ is in the connected phase as the mutual information between $B_{+}$ and $B_{-}$ is nonzero. However, as soon as the difference between $t_{a}$ and $t_{b}$, that is, $t_{a}-t_{b}$ equals $t_{Scr}^{(q)}$, the entanglement wedge associated with $B_{+}\cup B_{-}$ is in the disconnected phase as mutual information between $B_{+}$ and $B_{-}$ vanishes.
Finally, we will compute the Page time $(t_{P}^{(q)})$ for the QCSBH. This can be obtained by equating the leading order result of eq.(III) and eq.(39). The expression of $(t_{P}^{(q)})$ reads
@@
$t_{P}^{(q)}$$\approx$ $t_{P}^{(S)}\left(1-\frac{\tilde{\omega}G_{0}}{(r^{(S)})^{2}}\right)$ (44)
$=$ $t_{P}^{(S)}\left(1-\frac{\pi\tilde{\omega}}{S_{BH}^{(S)}}\right)$
where $t_{P}^{(S)}$ is the Page time for the standard Schwarzschild black hole, the expression of $t_{P}^{(S)}$ reads $t_{P}^{(S)}=\frac{3\beta^{S}}{\pi c}S_{BH}^{S}$. As we have aleady mention that the island starts contribute to the von-Neumann entropy of Hawking radiation just after the Page time. Therefore, our result suggests that for a quantum-corrected Schwarzschild black hole, island formation begins earlier compared to the standard Schwarzschild black hole. The fact that the island formation begins earlier in case of the QCSBH compared to the standard Schwarzschild black hole indicates that the island formation needed for getting the correct time evolution of von-Neumann entropy of Hawking radiation is favoured by the QCSBH which arises from the asymptotic safe theory of quantum gravity. The reason behind this can be understood from the following observation. It is clear that quantum gravity plays an important role in the resolution of the information loss paradox as the island formalism is an effective quantum gravity theory. The QCSBH solution, on the other hand, is derived from an effective action of gravity and captures some elements of quantum gravity which in turn helps in the island formation process earlier needed to get the correct time evolution of the von Neumann entropy of Hawking radiation.
@@
$=$ $\textrm{min}\leavevmode\nobreak\ \mathop{\textrm{ext}}_{\mathcal{\mathrm{I}}}\bigg{\{}\frac{\textrm{Area}(\partial I)}{4G_{N}}+S_{vN}(B_{+})+S_{vN}(B_{-})\bigg{\}}\leavevmode\nobreak\ .$
In the first line we make use of the fact that the state on the full Cauchy slice is a pure state and based upon this observation one can write $S_{vN}(B_{+}\cup B_{-})=S_{vN}(I\cup R)$. Furthermore, it is subject to the condition of vanishing of mutual information between $B_{+}$ and $B_{-}$. In the final form of the formula we have incorporated the the fact $t_{a}-t_{b}=|r^{*}(a)-r^{*}(b)|$ which in turn fixes one of the island parameters, namely, $t_{a}$ and only ``$a"$ is left to be extremized. We propose the above formula for evaluating the fine-grained entropy of Hawking radiation for an eternal black hole.
@@ III Before Page time scenario and the emergence of Hartman-Maldacena time
$S(R)$$\approx$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{\pi}\right)\sqrt{f(b)}\right]+\left(\frac{c}{6}\right)\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)^{2}\leavevmode\nobreak\ \leavevmode\nobreak\ .$
@@
$S(R)$$\approx$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{\pi}\right)\sqrt{f(b)}\right]+\left(\frac{c}{3}\right)\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ .$
From the above given expressions one can note that in the early time domain $S(R)$ has quadratic time dependence, that is $S(R)\sim t_{b}^{2}$. On the other hand in the late time domain, $S(R)$ grows linearly over observer’s time, $S(R)\sim t_{b}$ which matches with Hawking’s suggestion. This observation is agreed with Hartman and Maldacena (2013).
In this section our aim is to compute the mutual information between the matter fields localized on region $R_{+}$ and $R_{-}$. The reason behind this interest is the following. In the last section, we have already seen the importance of mutual information in obtaining the Page curve. In the before Page time scenario, although there is no island contribution but it will be very interesting to see how does the mutual information evolves over time. To compute the von Neumann entropy of matter fields on $R_{\pm}$, we will use the formula given below Calabrese and Cardy (2009)
@@
$d(b_{\pm},e_{\pm})$ $=$ $\sqrt{2F(b)F(e)e^{\kappa^{(q)}r^{*}(b)}[\cosh(\kappa^{(q)}r^{*}(b))-\cosh(\kappa^{(q)}t_{b})]}\leavevmode\nobreak\ .$
@@
$S_{vN}(R_{+})=S_{vN}(R_{-})=\left(\frac{c}{6}\right)\log\left[2\left(\frac{\beta^{(q)}}{2\pi}\right)^{2}\sqrt{f(b)f(e)}\left\{|\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)-\cosh\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)|\right\}\right]\leavevmode\nobreak\ .$ (55)
@@
$=$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(e)}\left\{\frac{|\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)-\cosh\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)|}{\cosh\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)}\right\}\right]\leavevmode\nobreak\ .$
For the sake of better understanding, we look at the behaviour of $I(R_{+}\cup R_{-})$ in two different time domains, namely in the early time and in the late time domain. In the early time domain, that is when $t_{b}\ll\beta^{(q)}$, the expression of the mutual information between $R_{+}$ and $R_{-}$ reduces to the following form
$I(R_{+}:R_{-})$$\approx$ $\left(\frac{c}{3}\right)\left[\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(e)}\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right]-\mathrm{sech}\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)-\left(\frac{2\pi^{2}}{\beta^{(q)^{2}}}\right)\left\{1+\mathrm{sech}\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right\}t_{b}^{2}\right]\leavevmode\nobreak\ .$
The above expression suggests that at the very begining, that is, at $t_{b}=0$, $I(R_{+}:R_{-})$ is non-zero and finite. It then decreases with respect to the observer’s time $t_{b}$. On the other hand at late time, that is for $t_{b}\gg\beta^{(q)}$, $I(R_{+}:R_{-})$ boils down to the following form
$I(R_{+}:R_{-})$$\approx$ $\left(\frac{c}{3}\right)\left[\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(e)}\right]-2\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)e^{-\left(\frac{2\pi t_{b}}{\beta^{(q)}}\right)}\right]\leavevmode\nobreak\ .$ (58)
The above result of $I(R_{+}:R_{-})$ implies that at late time $I(R_{+}:R_{-})$ increases with respect to $t_{b}$. This implies the behaviour of $I(R_{+}:R_{-})$ in two different time domains are different, in particular, at the early time $I(R_{+}:R_{-})$ is a decreasing function with respect to time and in the late time domain it is increasing. This in turn means that there is a lowest value of $I(R_{+}:R_{-})$ which characterizes this change in its nature. In order to have a concrete idea about the possible lowest value, we plot the general expression of $I(R_{+}:R_{-})$ (eq.(III)) with respect to the observer’s time $t_{b}$. In Fig.(4) we have given this where the vertical axis represents $\mathcal{I}_{R}=\left(\frac{3}{c}\right)I(R_{+}:R_{-})$ and the horizontal axis depicts $t_{b}$. Further, we have set $r^{(S)}=1$, $b=2r^{(S)}=2$ and $G_{0}=1$. We have also taken the limit $e\rightarrow\infty$. We observe that irrespective of the value for the quantum-correction paramater $\tilde{\omega}$, the mutual information between $R_{+}$ and $R_{-}$ vanishes for a particular value of the observer’s time $t_{b}$. This observation motivates us to look for a analytical expression corresponding to this particular time scale, namely, $t_{R}$.
Figure 4: The figure shows the variation of $\mathcal{I}_{R}$ with respect to the observer’s time $t_{b}$ for different values of $\tilde{\omega}$. The red, blue, and green curves represent the results for $\tilde{\omega}=0,0.1,0.2$ respectively. Here, we have set $r^{(S)}=1,b=2r^{(S)}=2$ and $G_{0}=1$.
The vanishing of the mutual information at $t_{b}=t_{R}$ means that the entanglement wedge associated with $R_{+}\cup R_{-}$ become disconnected. This was first discussed in Roy Chowdhury et al. (2022, 2023) with the help of the following proposal
@@
$t_{R}=$ $\left(\frac{\beta^{(q)}}{2\pi}\right)\cosh^{-1}\left\{\left(\frac{\frac{\beta^{(q)}}{2\pi}\sqrt{f(e)}}{1+\frac{\beta^{(q)}}{2\pi}\sqrt{f(e)}}\right)\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right\}.$
@@
$S_{vN}^{t_{b}=t_{R}}(R_{+}\cup R_{-})$ $=$ $\frac{c}{3}\log\left[\frac{\left(\frac{\beta^{(q)}\sqrt{f(e)}}{2\pi}\right)^{2}}{1+\frac{\beta^{(q)}\sqrt{f(e)}}{2\pi}}\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right]$
@@
So, our proposal suggests that in the time domain $0\leq t_{b}<t_{R}$, the mutual information between the matter fields on $R_{+}$ and $R_{-}$ is non zero and the entanglement wedge associated with $R_{+}\cup R_{-}$ is in its connected pahse. However, at $t_{b}=t_{R}$ the mutual correlation vanishes ($I(R_{+}\cup R_{-})=0$), this means that at $t_{b}=t_{R}$ the entanglement wedge corresponding to $R_{+}\cup R_{-}$ is in the disconnected pahse. These observations strongly indicate that this time-scale $t_{R}$ is nothing but the Hartman-Maldacena time $t_{HM}$ Hartman and Maldacena (2013). At the particular time-scale $t_{b}=t_{R}$, the entanglement surface corresponding to $R_{+}\cup R_{-}$ splits and leads to $S_{vN}(R_{+}\cup R_{-})=S_{vN}(R_{+})+S_{vN}(R_{-})$ which results in a vanishing mutual information between $R_{+}$ and $R_{-}$. This particular feature of $t_{R}$ suggests that it is precisely the Hartman-Maldacena time corresponding to the time-evolution of mutual information $I(R_{+}:R_{-})$ at which it saturates to the value $I(R_{+}:R_{-})=0$. This observation is in line with the one given in Grimaldi et al. (2022). Further, the expression of $I(R_{+}:R_{-})$ at $t_{b}=\beta^{(q)}$ (by substituting $t_{b}=\beta^{(q)}$ in eq.(III)), is given by
$I(R_{+}:R_{-})$ $=$ $\left(\frac{c}{3}\right)\log\left[\left(\frac{\beta^{(q)}}{2\pi}\right)\sqrt{f(e)}\cosh\left(\frac{2\pi r^{*}(b)}{\beta^{(q)}}\right)\right]\leavevmode\nobreak\ .$